PROOF OF ( X+1/X) > or = 2 , for a>0 and equality holds for a=1.

PROOF OF
( X+1/X) > or = 2   , for a>0 and
                                equality holds for a=1.

Grade:12th pass

2 Answers

Saurabh Koranglekar
5 years ago
using AM >= GM

( X+1/X)./2 >= (X*(1./X))^1/2

( X+1/X) >= 2*1

or
(X+1/X) >= 2
Deepak Kumar Shringi
5 years ago
useArithmetic Mean is always ≥Geometric Mean.

We can write:

(x+1/x )/2≥x⋅1/x

Thus,x+1/x more thsan equal to 2 but this is true when x>0
but if its less than zero
AM<GM
then x+1/x will be less than equal to 2